OpenStudy (meilendurcer):

How does finding the square root of a number compare to finding the cube root of a number? Use the number 64 in your explanation.

OpenStudy (anonymous):

square-root: \[x^2=64\\ x^2-64=0\\ (x-8)(x+8)=0\\ x-8=0\qquad x+8=0\\ x=8\qquad x=-8\] cube-root: \[x^3=64\\ x^3-64=0\\ (x-4)(x^2+4x+16)=0\\ x-4=0\qquad x^2+4x+16=0\\ x=4\qquad\qquad\text{no real roots here} \]